When the United States finally converts officially to metric units, the Celsius

temperature scale will replace the Fahrenheit scale for everyday use. As a

familiarization exercise, find the Celsius temperature corresponding to

a. Cool room (68°F)

b. A hot summer day (95°F) and

c. A cold winter day (5°F)


1
Expert's answer
2022-03-03T12:04:54-0500
"\rm{^\circ C}=\frac{5}{9}({^\circ F}-32)"

a.

"\rm 68{^\circ F}=\rm\frac{5}{9}(68-32)=20\rm{^\circ C}"

b.

"\rm 95{^\circ F}=\rm\frac{5}{9}(95-32)=35\rm{^\circ C}"

c.

"\rm 5{^\circ F}=\rm\frac{5}{9}(5-32)=-15\rm{^\circ C}"